a. The payment of the 19, 000 insurance premium for two years in advance was originally recorded as Prepaid Insurance. One year of the policy is now expired. b. All employees earn a total of P10,000 per day for a five-day week beginning on Monday and ending Friday. They were paid for the workweek ending December 26. They worked on Monday, December 29, Tuesday, December 30, and Wednesday, December 31. c. The supplies account had a balance of P4,480 on January 1. During the year, P11,000 of supplies were bought. A year-end inventory showed that P6,400 worth of supplies are still on hand. d. Equipment costing, P588,000 has a useful life of five years with an P80,000 salvage value at the end of five years. Record the depreciation for the year.

1
Expert's answer
2022-03-07T11:04:11-0500

"depreciation for the year=(588000-80000)\times\frac{1}{5}=101600"